OpenStudy (anonymous):

The sum of two consecutive integers is 59. Write an equation that models this situation and find the values of the two integers.

OpenStudy (anonymous):

n+n+1=59;n=29;n+1=30

OpenStudy (mendicant_bias):

Alright, so here's what you know about the situation, expressed in math: \[a+b=59,\]\[a = b-1\]

OpenStudy (mendicant_bias):

From here you can solve the system of equations. This is how you understand it, not just how you solve it.

OpenStudy (anonymous):

Thank You

OpenStudy (mendicant_bias):

No problem. Good luck.